# Numerical simulation of type II primordial black hole formation

### v5

@@ 4.1 PBH mass with the Gaussian initial fluctuation
[…]
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [4] already pointed out. However, it will correctly describe the accretion into the PBHs for sufficiently large times once we have a quasi-stationary flow into the PBH. The integrated solution is,
@@
where $M_{f}$ is the final BH mass. We fix the values of $M_{f}$ and $F$ by fitting Eq. (4.2) to the numerical data in a sufficiently late finite time interval denoted in Fig. 10 together with the time evolution of the mass for each value of $\mu$. The values of $F$ obtained by the fitting are listed in Table 2. The value of $F$ decreases with increasing $\mu$, which means the accretion is getting smaller. This behavior is consistent with the results in Ref. [45].
Figure 10: The time evolution of the BH mass for each value of $\mu$. The black dashed curves are fitting curves with Eq. (4.2) including the two parameters $M_{f}$ and $F$. The fitting region is taken for the interval between $t=450t_{H}$ and $t=500t_{H}$.
$\mu$ $M_{f}[M_{\text{H}}]$ $F$
@@
Table 2: The values of $F$ and $M_{f}$ for each amplitude of $\mu$ result from the fitting in Fig. 10.
Let us check the $\mu$ dependence of the PBH mass. In Fig. 11, we plot the initial and final PBH mass values for each value of $\mu$.
Figure 11: PBH mass in terms of the horizon mass at horizon entry $M_{H}$ to the amplitude of fluctuation $\mu$. In this case, in terms of fluctuation types, $\mu=1.2$ is classified as type I, while $\mu=1.4\>\text{to}\>2.2$ as type II, as seen in Fig. 2. In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type I, while $\mu=1.8$ to $2.2$ as type II, as seen in Figs. 6 and 9.
The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check the profile dependence in the next subsection.
@@ 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
In order to construct a specific initial profile that can be contrasted with the former Gaussian profile, let us try to utilize the three-zone profile as a reference, which is an extension of the model used in Ref. [1].
To realize a similar situation to the schematic conformal diagram, let us consider the closed FLRW region for the initial overdense region. Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
@@
where $k^{-1}$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region. Although this expression is relevant for the region $r\leq r_{a}$, we simply extend this expression in the region $r>r_{a}$, and to realize the condition $\zeta(L)=0$, we use the following specific functional form of $\zeta(r)$:
@@
where $\zeta_{\text{ref}}\coloneqq 0.001e^{-\frac{1}{2}k^{2}r^{2}}$ and $\tilde{\mu}\coloneqq k\chi_{a}/\pi$. The overall factor $\tilde{\mu}$ is multiplied for convenience to realize the background flat FLRW universe in the limit $\tilde{\mu}=0$. The value of $l$ is fixed as $k^{-1}=10H_{\rm b}^{-1}$ as before, and the value of $\chi_{a}$ is set as $\pi\tilde{\mu}/k$ for a given value of $\tilde{\mu}$.
@@
The initial profiles of $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs. 12 and 13. The tail profile is different from the original three-zone model, but it is similar to the three-zone profile around the peak of $\zeta$. This profile also exhibits the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
Figure 12: Initial three-zone profile of $\zeta(r)$ (left) and areal radius $R(r)$ (right) with respect to the amplitude parameter $\tilde{\mu}$. Taking the amplitude parameter $\tilde{\mu}$ larger, a type II fluctuation appeared for $\tilde{\mu}$ larger than $0.80$ with the neck $\partial_{r}R=0$. The original profile near the peak $\zeta_{\rm cFLRW}$ [1] is plotted in Fig. 19 in Appendix B.
Figure 13: Compaction function $\mathcal{C}_{\text{SS}}(r)$ for thnon-linearprofile. The horizontal axis is the coordinate radius with scale-up $z/L$ (left)/ the areal radius $R$ (right). By taking the amplitude $\tilde{\mu}$ larger, the number of peaks becomes two for $\tilde{\mu}$ = 0.89, corresponding to type II fluctuation.
@@
The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it. In contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig. 14, even though they are classified as type I fluctuations and formations for $\tilde{\mu}\leq 0.88$. This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases with the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation. We also find the gap between the parameter region of type I/II fluctuations and type I/II PBHs, as indicated in Fig. 14.
Figure 14: Initial PBH mass with respect to the amplitude of fluctuation for the three-zone profile.
@@
The decreasing behavior of the initial PBH mass with the amplitude can be easily predicted by focusing on the profile of the compaction function as a function of the areal radius. The initial PBH mass would be roughly estimated by the areal radius of the peak of the compaction function at the horizon entry. Therefore, if the areal radius around the peak of the compaction function increases and decreases with the amplitude of the perturbation profile, the initial PBH mass is expected to increase and decrease, respectively. In the case of the Gaussian profile, one can find that the areal radius around the peak of the compaction function increases with the value of the amplitude $\mu$ as the right panel of Fig. 2, while it decreases with $\tilde{\mu}$ in the case of the specific profile (4.4) as the right one of Fig. 13.
@@ 5 Conclusion
We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation. As in the dust case reported in KHW, the PBH formation process can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$). Specifically, if there is a bifurcating trapping horizon, the PBH is classified into type II, and type I otherwise. In the dust case, KHW reported that a type I fluctuation results in a type I PBH, and a type II fluctuation does in a type II PBH. However, in the case of radiation fluids, this is not true. We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of trapping horizons (formation type). That is, we found that a type II fluctuation may result in a type I PBH. Therefore, we propose classifying PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region. Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ initial hypersurface beyond the great circle. Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the initial time slice of $S^{3}$ in the overdensity region closes in the limit of large-density perturbation, which indicates the area of the edge is vanishing. However, this is not the case for the Gaussian-type initial fluctuation profile, while the other profile exhibits decreasing mass with increasing initial amplitude. Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile. Thus, systematic investigation of the profile dependence would be needed, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [19, 45]. In addition to the profile dependence, the dependence on the equation of state $p=w\rho$ is also an interesting issue, which fills the gap between KHW and this work [20].
@@ Acknowledgments
The authors are grateful to K.-I. Nakao, B.J. Carr, T. Igata, and J.M.M. Senovilla for useful comments and fruitful discussion. A.E. acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University). T.H. is supported in part by JSPS KAKENHI Grant Nos. 19K03876 and 20H05853. D.S. would like to take this opportunity to thank the “Nagoya University Interdisciplinary Frontier Fellowship” supported by Nagoya University and JST, the establishment of university fellowships towards the creation of science technology innovation, Grant Number JPMJFS2120. C.Y. is supported in part by JSPS KAKENHI Grant Nos. 20H05850 and 20H05853. C.Y. and T.H. acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
@@ Appendix A Conformal diagrams of the three-zone model for the dust case
In the three-zone model [7], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background. We also note that the intermediate underdense region must be introduced to compensate for the mass excess of the innermost region and connect the spacetime to the background spacetime without a mass gap. Without the intermediate underdense region, this model cannot be the exact solution, even for the dust case.
@@
For the underdense region with a perfect fluid with $w=1/3$, it would not be tractable to obtain any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. Despite this great simplification, the model provides a very powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [1]).
@@
The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig. 15, where future $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;*)$, past (+;)(+;*)( + ; ∗ ), and bifurcating $(0;*)$ trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively. For the dust case $w=0$, the BH apparent horizon and cosmological apparent horizon become future-outer (;)(-;-)( - ; - ) and past-inner (+;+)(+;+)( + ; + ), denoted by red -solid and blue-dashed lines, respectively.
@@
First, let us construct the spacetime diagram of the case of the type I fluctuation. The extracted regions for the patchwork in the type I case are described by the yellow shaded regions in Fig. 15, and the resultant spacetime diagram is shown in the left panel of Fig. 17. These conformal diagrams define the right and left-directed future null vectors as $l_{\rm out}$ and $l_{\rm in}$, respectively. For the closed FLRW region, on the boundary of the extracted region, we can find $\theta_{\rm out}>0$ and $\theta_{\rm in}<0$ at the maximum expansion. Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta_{\rm out}$ and $\theta_{\rm in}$. Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
@@
The extracted regions are yellow-shaded for the type II fluctuation in Fig. 16. Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion. Then we find $\theta_{\rm out}<0$ and $\theta_{\rm in}>0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram. Thus, the resultant conformal diagram for type II is given by the right panel of Fig. 17.
@@
Figure 15: Conformal diagrams to use for a patchwork for the spacetime of the type I PBH schematically in Fig. 17 for dust case $w=0$. Dashed gray curves represent $R$-levels. Red(blue)-shaded regions represent future(past) trapped regions. Future-outer $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;-)$, past-outer (+;)(+;-)( + ; - ), future-inner (;+)(-;+)( - ; + ), and past-inner (+;+)(+;+)( + ; + ) trapping horizons are denoted by red-solid, blue -solid, red -dashed, and blue -dashed lines, respectively.
@@
Figure 16: Same as Fig. 15 but for the type II PBH. We use the yellow-shaded regions for a patchwork.
(a) Type I PBH
(b) Type II PBH
Figure 17: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons. The left figure defines a type I (type A) PBH as a typical case PBH causal structure. The right one is constructed by changing the cutting diagram range and defining a type II (type B) PBH.
@@
Figure 18: Conformal diagrams of the closed and flat FLRW solutions for the radiation fluid case $w=1/3$. In this case, all trapping horizons are null. Future-degenerate $(\theta_{-};\mathcal{L}_{-}\theta_{+})|_{\theta_{+}=0}=(-;0)$ and past-degenerate $(+;0)$ trapping horizons are denoted by red-dotted and blue-dotted lines, respectively.

### v6

@@ 4.1 PBH mass with the Gaussian initial fluctuation
Let us introduce the mass of the PBH in two different ways at each characteristic time. The first is to calculate the Misner-Sharp mass at the apparent horizon at its first appearance with the future-trapped region (see Fig. 6). We refer to the mass at this time as the initial BH mass. In a type A PBH, the BH apparent horizon appears as a future-outer trapping horizon, while it appears as a bifurcating outer trapping horizon in type B. We also calculate the final PBH mass by estimating the effect of the accretion. The accretion process is well described by the mass accretion model, often called Novikov-Zeldovich or Bondi accretion [Zeldovich:1967lct, 1952MNRAS.112..195B]. In this accretion model, the time evolution of the mass is given by
@@
where $\rho_{b}=\rho_{b,0}(t_{0}/t)^{2}$ is the background energy density, and $F$ is a nondimensional constant that corresponds to the efficiency of accretion, which is commonly numerically found to be of order $\mathcal{O}(1)$ and will depend on the details of the accretion process. It is important to note that Eq. (4.1) is not valid at the time of formation of the trapping horizon since it neglects the cosmological expansion as Carr and Hawking (1974) [10.1093/mnras/168.2.399] already pointed out. However, once we have a quasi-stationary flow into the PBH, it will correctly describe the accretion into the PBHs for sufficiently long times. The integrated solution is,
@@
where $M_{f}$ is the final BH mass. We find the values of $M_{f}$ and $F$ by fitting Eq. (4.2) to the numerical data in a sufficiently late finite time interval denoted in Fig. 11 together with the time evolution of the mass for each value of $\mu$. The values of $F$ obtained by the fitting are listed in Table 2. The value of $F$ decreases with increasing $\mu$, which means the accretion is getting smaller. This behavior is consistent with the results in Ref. [escriva_2021].
Figure 11: The time evolution of the BH mass for $\zeta(r)$ given by Eq. (2.5) for $\mu=1.2,1.4,1.6,1.8,2.0$ and $2.2$. The black dashed curves are fitting curves with Eq. (4.2) including the two parameters $M_{f}$ and $F$. The fitting region is taken for the interval between $t=450t_{H}$ and $t=500t_{H}$. The values of fitting parameters are shown in Table 2.
$\mu$ $M_{f}/M_{\text{H}}$ $F$
@@
Table 2: The values of $F$ and $M_{f}$ for $\zeta(r)$ given by Eq. (2.5) for $\mu=1.2,1.4,1.6,1.8,2.0$ and $2.2$ obtained from the fitting in Fig. 11.
Let us check the $\mu$ dependence of the PBH mass. In Fig. 12, we plot the initial and final PBH mass values for each value of $\mu$. This behavior can be seen for an amplitude much larger than the critical regime, where $M$ obeys a power law against $(\mu-\mu_{c})$ [PhysRevLett.70.9, Evans:1994pj, Koike:1995jm, Niemeyer:1999ak, Musco:2012au].
Figure 12: PBH mass in terms of the horizon mass at horizon entry $M_{H}$ to the amplitude of fluctuation $\mu$ for $\zeta(r)$ given by Eq. (2.5). In this case, in terms of fluctuation types, $\mu=1.2$ is classified as type I, while $\mu=1.4\>\text{to}\>2.2$ as type II, as seen in Fig. 2. In terms of formation types, on the other hand, $\mu=1.2$ to $1.6$ are classified as type A, while $\mu=1.8$ to $2.2$ as type B, as seen in Figs. 6 and 10. We can see that the initial and final PBH masses get larger with increasing amplitude $\mu$.
The initial and final PBH masses get larger with increasing amplitude $\mu$. We will check whether this behavior appears for a different profile or not in the next subsection.
@@ 4.2 PBH mass with the initial fluctuation motivated by the three-zone model
[…]
To construct a specific initial profile that can be compared with the former Gaussian profile as Eq. (2.5) and Fig. 1, let us try to utilize the three-zone profile as a reference, an extension of the model used in Ref. [PhysRevD.83.124025]. Let us consider the closed FLRW region for the initial overdense region to realize a situation similar to the schematic conformal diagram. Then, as is shown in the Appendix B, under the boundary condition $\zeta=0$ at the surface of the overdense region, we obtain the following form of $\zeta(r)$:
@@
where $k^{-1}$ is the comoving radius of $S^{3}$ associated with the closed FLRW and $0<\chi_{a}\leq\pi/k$ describes the radial coordinate of the edge of the closed FLRW region. Although this expression is relevant for the region $r\leq r_{a}$, this functional form cannot be directly used as the initial profile for our numerical simulation. This is because, in our simulation, we suppose that the initial profile is continuously connected to the flat FLRW background, namely, $\zeta(r)$ should vanish for $r>L$. To realize this condition, first, we simply extend the expression (4.3) to the region $r>r_{a}$, and use the following specific functional form of $\zeta(r)$ for the initial condition of our simulation:
@@
where $\zeta_{\text{ref}}\coloneqq 0.001e^{-\frac{1}{2}k^{2}r^{2}}$ and $\tilde{\mu}\coloneqq k\chi_{a}/\pi$. The overall factor $\tilde{\mu}$ is multiplied for convenience to realize the background flat FLRW universe in the limit $\tilde{\mu}=0$. The value of $l$ is fixed as $k^{-1}=10H_{\rm b}^{-1}$ as before, and the value of $\chi_{a}$ is set as $\pi\tilde{\mu}/k$ for a given value of $\tilde{\mu}$. This profile looks complicated and differs from the original three-zone model profile but is useful for showing a specific example of the non-monotonic behavior of the PBH mass.
@@
The initial profiles of $\zeta$, $R$ and the compaction function $\mathcal{C}_{\rm SS}$ are shown in Figs. 13 and 14. The tail profile is different from the original three-zone model, but it is similar to the three-zone profile near the peak of $\zeta$. This profile can also exhibit the type II feature with the neck structure for a large amplitude $(\tilde{\mu}\gtrsim 0.85)$.
Figure 13: Same as Fig. 1 but for $\zeta(r)$ given by Eq. (4.4). The functional form (4.4) is inspired by the three-zone model but modified for numerical simulations and is different from the Gaussian-shaped profile (2.5). We plot $\zeta(r)$ for the set of the amplitude parameter values $\tilde{\mu}=0.20,0.40,0.60,0.80,0.90$ and $0.92$. We can see that $R(r)$ is a monotonically increasing function for $\tilde{\mu}=0.20,0.40,0.60$ and $0.80$, while it increases, decreases and increases again as $r$ increases for $\tilde{\mu}=0.90$ and $0.92$. In the latter cases, $R(r)$ has a local maximum and a minimum with $\partial_{r}R=0$, where the minimum corresponds to the neck.
Figure 14: Same as Fig. 2 but for $\zeta(r)$ given by Eq. (4.4). We can see that $\mathcal{C}_{\text{SS}}(r)$ takes the only one maximum, which is smaller than $1/2$, for $\tilde{\mu}=0.20,0.40,0.60$, and $0.80$, while it takes two distinct maxima, which are equal to $1/2$, and a single minimum between the two maxima for $\tilde{\mu}=0.90$ and $0.92$. The latter cases correspond to type II fluctuations as shown in Appendix C. We can see that $\mathcal{C}_{\text{SS}}(r)$ takes the only maximum or the two maxima at the smaller areal radius $R$ for the larger values of $\tilde{\mu}$. This implies that the mass of the formed black hole decreases as $\tilde{\mu}$ increases as seen in Fig. 15. This is not the case for the Gaussian-shaped functional forms for $\zeta(r)$ shown in Fig. 2, for which the mass of the formed black hole monotonically increases as $\tilde{\mu}$ increases as shown in Fig. 12 at least for the parameters computed here.
@@
The behavior of the horizon configuration is similar to the case of the Gaussian profile, and we do not explicitly show it. However, in contrast to the Gaussian profile, the initial PBH mass formed by this profile decreases as the amplitude of fluctuation $\tilde{\mu}$ increases for $\tilde{\mu}>0.65$ as shown in Fig. 15, even though they are classified as type I fluctuations and type A formations for $\tilde{\mu}\leq 0.88$. Again, this is well beyond the near-critical regime. This result, as well as the behavior of the Gaussian profile, indicates that whether the initial PBH mass increases or decreases as a function of the amplitude parameter is not only up to its type but also strongly dependent on its profile, together with the ambiguity in quantifying the amplitude of non-linear perturbation. We also find the gap between the parameter region of type I/II and type A/B PBHs, as indicated in Fig. 15.
Figure 15: Initial PBH mass with respect to the amplitude of fluctuation for $\zeta(r)$ given by Eq. (4.4) We can see that the initial PBH mass decreases as the amplitude parameter $\tilde{\mu}$ is increased for $\tilde{\mu}\gtrsim 0.65$, after it increases for $\tilde{\mu}_{c}<\tilde{\mu}\lesssim 0.65$.
@@
The decreasing behavior of the initial PBH mass with the amplitude can be easily predicted by focusing on the profile of the compaction function as a function of the areal radius. The initial PBH mass would be roughly estimated by the areal radius at the peak of the compaction function at the horizon entry. Therefore, if the areal radius at the peak of the compaction function increases and decreases with increasing the amplitude of the perturbation profile, the initial PBH mass is expected to increase and decrease as well, respectively. In the case of the Gaussian profile (2.5), one can find that the areal radius at the peak of the compaction function increases with increasing the value of the amplitude $\mu$ as seen in the right panel of Fig. 2, while it decreases with increasing $\tilde{\mu}$ in the case of the specific profile (4.4) as seen in the right one of Fig. 14.
@@ 5 Conclusion
We have numerically investigated the PBH formation from a type II initial fluctuation $\zeta(r)$, which is defined by the non-monotonicity of the areal radius $R(r)\propto re^{\zeta}$ with respect to the coordinate radius $r$ on the spatial geometry, i.e., the existence of $\partial_{r}R(r)=0$ (called neck or throat structure), in the radiation-dominated situation. As in the dust case reported in KHW, the initial fluctuation can be classified into types I and II according to the configuration of the trapped regions $R(t,r)\leq 2M(t,r)$ (or trapping horizons $R(t,r)=2M(t,r)$). Specifically, if there is a bifurcating trapping horizon, the PBH is classified as type B, and otherwise, it is classified as type A. In the dust case, KHW reported that a type I fluctuation results in a type A PBH and a type II fluctuation in a type B PBH. However, in the case of radiation fluids, this is not true. We found a gap between the classifications regarding the shape of the initial spatial geometry (fluctuation type) and the structure of the trapping horizons (formation type). We found that a type II fluctuation may result in a type A PBH, which we will call type II-A PBH. Therefore, we propose classifying type A/B PBHs based on the existence of the bifurcating trapping horizon independently of the type of the initial fluctuation.
In the leading order of the long-wavelength limit, the overdense region can be roughly approximated by the closed FLRW universe, and the amplitude of the density perturbation can be translated into the value of the radial coordinate at the edge of the closed FLRW region. Then, the type II fluctuation indicates that the edge is located in the other hemisphere of $S^{3}$ beyond the great circle on the initial hypersurface. Therefore, in the case of type II PBHs, one may expect that the PBH mass would decrease with increasing the amplitude of initial fluctuation since the edge of the overdense region in $S^{3}$ closes in the limit of large-density perturbation, which indicates that the area of the edge is vanishing. However, this is not the case for the Gaussian-type initial fluctuation profile, whereas the other profile exhibits decreasing mass with increasing initial amplitude. Therefore, we conclude that the behavior of the initial PBH mass depends on not only the type of fluctuation but also the fluctuation profile. Thus, a more systematic investigation would be needed to reveal the details of the profile dependence, e.g., by using the shape parameter defined as the curvature peak of the compaction function $\mathcal{C}_{\text{SS}}$ [Escriva:2019phb, escriva_2021]. In addition to profile dependence, dependence on the equation of state $p=w\rho$ is also an interesting issue, filling the gap between KHW and this work [escriva_2021analy].
@@ Acknowledgments
The authors are grateful to I. Musco, K.-I. Nakao, B.J. Carr, T. Igata, and J.M.M. Senovilla for useful comments and fruitful discussion. A.E. acknowledges support from the JSPS Postdoctoral Fellowships for Research in Japan (Graduate School of Sciences, Nagoya University). The authors are supported in part by the JSPS KAKENHI Grant Nos. 19K03876(TH), 20H05850(CY), 20H05853(TH, CY), and 24KJ1223(DS). KU would like to take this opportunity to thank the “THERS Make New Standards Program for the Next Generation Researchers” supported by JST SPRING, Grant Number JPMJSPS2125. CY and TH acknowledge the hospitality at APCTP during the focus research program ”Black Hole and Gravitational Waves: from modified theories of gravity to data analysis”, where part of this work was discussed.
@@ Appendix A Conformal diagrams of the three-zone model for the dust case
In the three-zone model [PhysRevD.88.084051], the spacetime of PBH is described by the patchwork of the three different regions: the innermost overdense region is given by a spatially closed FLRW universe, the intermediate underdense region, and the outermost flat FLRW universe as the background. We also note that the intermediate underdense region must be introduced to compensate for the innermost region’s mass excess and connect the spacetime to the background spacetime without a mass gap. Without the intermediate underdense region, this model cannot be an exact solution, even for the dust case.
@@
For the underdense region with a perfect fluid with $w=1/3$, obtaining any analytical and exact solutions of such dynamical inhomogeneous spacetimes with non-linear sound waves would not be tractable. Therefore, we stick to the three-zone model with dust. Furthermore, for simplicity, we focus on the model where the underdense region is described by the Schwarzschild solution, where the matching is possible without any singular hypersurface for the dust case. Despite this significant simplification, the model provides a powerful tool for capturing the overall dynamics of PBH and the structure of trapping horizons. Note that one can also approach the same problem by using the Lemaître-Tolman-Bondi exact dust solution (c.f. [PhysRevD.83.124025]).
@@
The conformal diagrams for the three spacetimes of which the three-zone model consists are shown in Fig. 16, where future, past, and bifurcating trapping horizons are denoted by red lines, blue lines, and the intersection point between them in the Schwarzschild region, respectively. For the dust case $w=0$, the apparent horizon and cosmological apparent horizon become future-outer and past-inner, denoted by red-solid and blue-dashed lines, respectively. See Table 1 for the classification of trapping horiozons.
@@
First, let us construct the spacetime diagram of the case of the type I fluctuation. The extracted regions for the patchwork in the type I (type A) case are described by the yellow shaded regions in Fig. 16, and the resultant spacetime diagram is shown in the left panel of Fig. 18. For the closed FLRW region, on the boundary of the extracted region, we can find $\theta<0$ and $\theta^{\prime}>0$ at the maximum expansion. Therefore, in the outside Schwarzschild region, we also require the same sign of $\theta$ and $\theta^{\prime}$. Similarly, one can understand the matching of the Schwarzsdhild region and the background spatially flat FLRW spacetime.
@@
The extracted regions are yellow-shaded for the type II fluctuation in Fig. 17. Since the type II fluctuation should have a maximal value of the areal radius, the boundary of the closed FLRW region must be outside the central point at which the areal radius takes the maximum value corresponding to the area of the big sphere at the maximum expansion. Then we find $\theta>0$ and $\theta^{\prime}<0$ on the boundary, which can be matched with the Schwarzschild region only in the left part of the Schwarzschild conformal diagram. Thus, the resultant conformal diagram for type II (type B) is given by the right panel of Fig. 18.
For reference, Fig. 19 gives the conformal diagrams of the closed and flat FLRW spacetimes with $w=1/3$ in the left and right panels, respectively.
@@
Figure 16: Conformal diagrams to use for a patchwork for the spacetime of the type I (type A) PBH schematically in Fig. 18 for dust case $w=0$. The left, middle, and right panels show the conformal diagrams of the closed FLRW, the Schwarzschild, and the flat FLRW spacetimes, respectively. Dashed gray curves represent $R$-levels. Red (blue)-shaded regions represent future (past) trapped regions. Future-outer, past-outer, future-inner, and past-inner trapping horizons are denoted by red-solid, blue-solid, red-dashed, and blue-dashed lines, respectively. See Table 1 for the classification of trapping horizons.
@@
Figure 17: Same as Fig. 16 but for the type II (type B) PBH. We use the yellow-shaded regions for a patchwork.
(a) Type A PBH
(b) Type B PBH
Figure 18: Conformal diagrams of PBH for dust case $w=0$ as a patchwork of the three diagrams and trapping horizons. The left panel defines a type I (type A) PBH as a typical case PBH causal structure. The right one is that for a type II (type B) PBH.
@@
Figure 19: Conformal diagrams of the closed and flat FLRW solutions in the left and right panels, respectively, for the radiation fluid case $w=1/3$. In this case, all trapping horizons are null. Future-degenerate and past-degenerate trapping horizons are denoted by red-dotted and blue-dotted lines, respectively. See Table 1 for the classification of trapping horizons.